OpenStudy (anonymous):

Is what i am doing correct Question : 2log(7) - log 14

OpenStudy (anonymous):

What are you trying to do?

OpenStudy (anonymous):

log(49) - log(14) log(49/14)

OpenStudy (anonymous):

That looks right

OpenStudy (shiraz14):

@╰☆╮Openstudier╰☆╮ : Yes, you're correct.

OpenStudy (anonymous):

Since 2log7 = log 7^2

OpenStudy (campbell_st):

well if you simply the fraction 49/14 = 7/2 so its log(7/2) = log(7) - log(2)

OpenStudy (anonymous):

the answer given is log5(49)(sqrt 3)

OpenStudy (anonymous):

It depends if you need to simplify

OpenStudy (campbell_st):

so what is the base of the logs...?

OpenStudy (anonymous):

5

OpenStudy (campbell_st):

ok... so they are both the same base...?

OpenStudy (anonymous):

So pull the base out front and get 5log

OpenStudy (anonymous):

we pull out the power right

OpenStudy (campbell_st):

what is the whole question... in what you have posted the question doesn't match the answer...

OpenStudy (anonymous):

Question : 2log(7) - log 14

OpenStudy (anonymous):

What is the base?

OpenStudy (anonymous):

Simplify I assume, and the base is 10?

OpenStudy (anonymous):

simplify yes

OpenStudy (anonymous):

Ok the first part is log 7^2

OpenStudy (anonymous):

log(49) - log(14) log(49/14)

OpenStudy (campbell_st):

well the base doesn't matter as both logs are the same base.... so the normal log laws apply so looking at it and applying the index laws for logs \[2\log(7) - \log(14) = \log(7^2) - \log(14) = \log(49) - \log (14)\] now the division law for logs \[\log(49) - \log(14) = \log(\frac{49}{14}) = \log(\frac{7}{2})\] that the job done...

OpenStudy (anonymous):

If you are subtracting logs, then you divide them

OpenStudy (anonymous):

Like @campbell_st mentioned it's log(7/2)

OpenStudy (anonymous):

or log(7)-log(2) both work..

OpenStudy (anonymous):

I got that my point is the answer given is log5(49)(sqrt 3 )

OpenStudy (anonymous):

that**

OpenStudy (campbell_st):

well I can't see the connection between the question and answer....

OpenStudy (anonymous):

I don't see how the book got that, unless you were given a different base other than 10?

OpenStudy (anonymous):

WELL the answer is incoorrect then

OpenStudy (anonymous):

incorrect*

OpenStudy (ranga):

Can you post a screenshot/pic of the entire question?

OpenStudy (campbell_st):

based on the question and information you have provided.... unless there is something that been ommited...

OpenStudy (anonymous):

log(7/2) is given in the next exercise to the excersise which i am doing currently so the answer is clearly misprinted log(7/2) is correct

OpenStudy (anonymous):

THANK YOU ALL OF YOU

OpenStudy (campbell_st):

lol.... never trust the answer in the back of the book